OpenStudy (ujjwal):

A constant force acts on a particle and its displacement (y) is related to time (t) as t=sqrt(y) +3 When the velocity is zero, its displacement (in cm) is: a)0 b)1 c)2 d)3

OpenStudy (experimentx):

y(t) = (t-3)^2

OpenStudy (callisto):

I'm not sure if this is correct. Since velocity = displacement/time 0 = displacement/time displacement = 0

OpenStudy (anonymous):

DIFFERENTIATE IT WITH RESPECT TO TIME AND YOU WILL GET \[Y=\frac{\frac{dy}{dt}}{4}\] since give velocity is zero then displacement should also be zero.

OpenStudy (vincent-lyon.fr):

@ujjwal @Callisto velocity is not displacement over time but derivative of displacement wrt time. This is not the same, except in case of steady motion.

OpenStudy (ujjwal):

@ramkrishna how do you get that value of y.. here, dy/dt =4t+6.. isnt it?

OpenStudy (vincent-lyon.fr):

If y = (t-3)², then v = 2 (t-3) v is zero for t=3, so when v=0, then y=0.

OpenStudy (callisto):

I.... am sorry!!!!!!!!

OpenStudy (ujjwal):

I think the same way @Vincent-Lyon.Fr ..

OpenStudy (ujjwal):

@Callisto you needn't say sorry for that.. you tried to help and thanks a lot for that..

OpenStudy (vincent-lyon.fr):

@Callisto No problem. These are subtleties and lots of problems are solved with displacement /time.

OpenStudy (anonymous):

\[y=\sqrt{y}+3\] \[t-3=\sqrt{y}\] \[(t-3)*(t-3)=y\] \[v=2*(t-3)\] when v=0 t=3 using t=3 in the equation (t−3)∗(t−3)=y we get y=0

OpenStudy (ujjwal):

ok, thanks everybody!

OpenStudy (callisto):

And thanks everyone for correcting my mistakes :)